# The Unbounded Denominators Conjecture

### v3

@@ 7.4.1. A brief survey of the literature on Mason’s conjecture
Originally Kohnen and Mason [KM08, KM12] focused on the particular case (GMF) that the representation $\rho$ is monomial (induced from a one-dimensional character on a finite index subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$). They used the Rankin–Selberg method to prove the conjecture in the case of a generalized modular function (weight 0) without any zeros or poles on the extended upper-half plane [KM08, Theorem 1]. In fact Selberg’s paper [Sel65] that they used here had already considered vector-valued modular forms for the purpose of extending the Rankin–Selberg estimate into the noncongruence case. Kohnen and Mason [KM08, Theorem 2], again based on the Rankin–Selberg $L$-function method but now with a finer input from the Eichler–Shimura–Weil bound on Fourier coefficients of congruence cusp forms in weight $2$, also proved that when $\rho$ is induced from a linear character of a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$, the same result on generalized modular function units also holds if the condition on integer coefficients is relaxed to $S$-integer coefficients: a case that goes beneath the scope of our results here.
@@
In a sequel work [KM12], Kohnen and Mason used the Knopp–Mason canonical factorization [KM09] $f=f_{0}f_{1}$ (over $\mathbf{C}$) of a parabolic generalized modular function $f$ on a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$, where $f_{0}$ is a parabolic generalized modular function of a unitary character $\chi$, while $f_{1}$ is a parabolic generalized modular function without zeros or poles on the extended upper-half plane [KM03a]. Combining to their earlier method from [KM08], they thus proved that the unbounded denominators conjecture for the case of parabolic GMF is equivalent to the algebraicity of the first “few” Fourier coefficients of the component $f_{1}$ in the canonical factorization of $f$. As an application they proved Mason’s unbounded denominators conjecture for the case of a cuspidal parabolic GMF of weight 0 on a congruence group.
@@
In the case $n=2$ of two-dimensional representations, Mason’s conjecture was settled by Franc and Mason [Mas12, FM14], and extended further by Franc, Gannon and Mason [FGM18] to the stronger sense of only requiring the $p$-adic boundedness of the coefficients for a full density set of primes $p$. Their proof relies on the special incidence that the rank-$2$ local systems on $\mathbf{P}^{1}\smallsetminus\{0,1,\infty\}$ reduce to the Gauss hypergeometric equation, and the classical theory of hypergeometric functions. It is conceivable that the algebraicity part (over $\mathbf{Q}(x)$, respectively over the ring of classical modular forms) in Theorems 7.2.1 and 7.3.3 could likewise hold under a similar loosening of the integrality condition; but our proof does not yield to this. On the other hand, for representations of dimension $n\geq 3$, it is plain that the congruence property ceases to hold as in [FM14] if we relax $\mathbf{Z}\llbracket q\rrbracket$ to $\mathbf{Z}[1/S]\llbracket q\rrbracket$. The hypergeometric method was extended to three-dimensional representations ($n=3$) of $\mathrm{SL}_{2}(\mathbf{Z})$ by Franc–Mason [FM16a] and Marks [Mar15], and employed back in [FM16b] to derive certain cases of the original unbounded denominators conjecture.
@@ 7.4.2. Logarithmic vector-valued modular forms
If one drops the semisimplicity stipulation on $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ in the definition of a vector-valued modular form, the resulting structure has been named a logarithmic vector-valued modular form by Knopp and Mason [KM11]. They also do arise in conformal field theories, termed logarithmic (in place of rational). See, for example, Fuchs–Schweigert [FS19]. But now by Remark 7.2.3 the components of a weight zero vector-valued modular form with bounded denominators can certainly be transcendental over $\mathbf{C}(\lambda)$. Still the examples there are classical (congruence) modular forms, except of a higher weight. In Question 7.4.3 below we give an extension of the unbounded denominators problem over to the logarithmic setting. It remains outside the scope of our method as far as we could see. Before stating this question, we recall some basic facts concerning quasi-modular forms. Recall that the ring of quasi-modular forms $\widetilde{M}(\Gamma)$ for $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ may be identified with the ring generated by $E_{2}$ over the ring of classical holomorphic modular forms $M(\Gamma)$ of integral weight for $\Gamma$ [Zag08, Prop 20(ii)], and by [Zag08, Prop 20(i)] it is stable under the operator
@@
The ring of weakly holomorphic quasi-modular forms $\widetilde{M}^{!}(\Gamma)$ for $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ is the ring $\widetilde{M}(\Gamma)[1/\Delta]$.
@@
Question 7.4.3.
If a component $F_{j}(\tau)$ of a logarithmic vector-valued modular form for $\mathrm{SL}_{2}(\mathbf{Z})$ has a $\mathbf{Z}\llbracket q\rrbracket$ Fourier expansion, does $F_{j}(\tau)$ belong to the ring of weakly holomorphic quasi-modular forms $\widetilde{M}^{!}(\Gamma)$ for some congruence subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$?
@@
We now turn to some basic examples hinting towards a positive answer to question 7.4.3. Complementing Example 7.2.3 is the $\lambda$-pullback of the complete elliptic integral of the second kind:
@@
clearly a component of a logarithmic vector-valued modular form on $\mathrm{SL}_{2}(\mathbf{Z})$, whose $q$-expansion is in $\mathbf{Z}\llbracket q\rrbracket$. But one can indeed verify that
$\frac{2}{\pi}E(\lambda(q))=\frac{\vartheta_{3}\vartheta^{4}_{4}+4\theta\vartheta_{3}}{\vartheta^{3}_{3}}.$
@@
exhibiting this $\mathbf{Z}\llbracket q\rrbracket$ power series as an element of $\widetilde{M}^{!}(\Gamma)$ for the congruence subgroup $\Gamma=\Gamma(12)$, in accordance with Question 7.4.3.
@@ 7.4.4. Beyond $\mathrm{SL}_{2}(\mathbf{Z})$
Secondly, the mapping class groups $\Gamma_{g,n}=\mathrm{Mod}(S_{g,n})$ in signatures $(g,n)$ other than $(1,1),(1,0)$ or $(0,4)$ that we have implicitly been limiting to. Recall that $\Gamma_{1,1}\cong\Gamma_{1,0}=\mathrm{Mod}(\mathbf{T}^{2})=\mathrm{SL}_{2}(\mathbf{Z})$ and $\Gamma_{0,4}\cong\mathrm{PSL}_{2}(\mathbf{Z})\ltimes(\mathbf{Z}/2\times\mathbf{Z}/2)$, and correspondingly the discussion in the rational conformal field theory under §7.1 has been for the $1$-loop partition function with a complex torus ($g=1$) as the worldsheet [Gan06]. In a more recent research stream in two-dimensional conformal field theory, a higher genus extension of Zhu’s modularity theorem was recently obtained by Codogni [Cod20], on associating to any holomorphic vertex operator algebra a Teichmüller modular form in every signature $(g,n)$: a section of a tensor power $\lambda^{\otimes(c/2)}$ of the Hodge bundle over $\overline{\mathcal{M}_{g,n}}$, where the (doubled) weight $c$ is the central charge of the vertex algebra. This Teichmüller modular form is, up to the $c$-th power of a certain higher genus generalization [MT06] of the Dedekind eta function, equal to the partition function of the conformal field theory associated to the vertex algebra. At the very least, one could ask about extending the cruder algebraicity proviso of our Theorem 7.3.3 over to the more general setting of a component of a vector-valued Teichmüller modular form that has the appropriate integrality property.
@@ 7.4.5. Algebraic fundamental groups
Let $N\in\mathbf{N}$, let $K/\mathbf{Q}(\mu_{N})$ be a finite extension, and let $\pi:\mathcal{X}(N)\to\mathrm{Spec}\,{O_{K}}$ (“connected Néron model”) be the connected component containing the cusp $\infty$ in the smooth part of the minimal regular model of $X(N)$ over $\mathrm{Spec}\,{O_{K}}$. Thus the cusp $\infty$ extends to a morphism $\varepsilon:\mathrm{Spec}\,{O_{K}}\to\mathcal{X}(N)$.
@@
This follows rather formally by the argument of [Iha94, § 4 on page 252] and [Iha94, proof of Theorem 1 loc.cit. on pages 248–249], upon replacing Ihara’s function field $k(t)$ by the modular function field $K(X(N))$ and Ihara’s formal power series ring $\mathfrak{O}\llbracket t\rrbracket$ by $O_{K}\llbracket\lambda(\tau/N)/16\rrbracket$, taking account of Remark 6.3.1, and on using our Theorem 1.0.1 in place of Harbater’s arithmetic rationality input [Iha94, Claim 1A on page 248]. ∎
@@
by the Lerch–Chowla–Selberg formula making Bost’s capacitary condition [Bos99, Corollary 1.3] apply, and this is the isolated minimum value of the Faltings height across all elliptic curves. In practice this means that this complex torus has a “large” univalent complex-analytic uniformization (in the sense of conformal size from the origin $\infty$] and potential theory), sufficient to place this particular case of Theorem 7.4.4 to within the framework of arithmetic rationality — as opposed to algebraicity or holonomicity — theorems [Bos99, BCL09] on the algebraic curve $X(N)$. Can such an approach be continued to all $N$?
@@ 8. Acknowledgments
We would like to thank Yves André for a number of insightful remarks on the first version of this manuscript, leading in particular to § 2.4.1 as an alternative and simplified approach to Theorem 2.0.1. We would also like to thank Michael Barz, Jean-Benoît Bost, François Charles, Pierre Deligne, Cameron Franc, Igor Frenkel, Javier Fresán, Kenz Kallal, Mark Kisin, Geoffrey Mason, Alex Smith, Richard Taylor, John Voight, and Wadim Zudilin for useful remarks, suggestions, and corrections.

### v4

@@ 7.4.1. A brief survey of the literature on Mason’s conjecture
Originally Kohnen and Mason [KM08, KM12] focused on the particular case (generalized modular forms) that the representation $\rho$ is monomial. They used the Rankin–Selberg method to prove the conjecture in the case of a generalized modular function (weight 0) without any zeros or poles on the extended upper-half plane [KM08, Theorem 1]. In fact Selberg’s paper [Sel65] that they used here had already considered vector-valued modular forms for the purpose of extending the Rankin–Selberg estimate into the noncongruence case (see also § 7.4.8 below). Kohnen and Mason [KM08, Theorem 2], again based on the Rankin–Selberg $L$-function method but now with a finer input from the Eichler–Shimura–Weil bound on Fourier coefficients of congruence cusp forms in weight $2$, also proved that when $\rho$ is induced from a linear character of a congruence subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$, the same result on generalized modular function units also holds if the condition on integer coefficients is relaxed to $S$-integer coefficients: a case that goes beneath the scope of our results here.
@@
In a sequel work [KM12], Kohnen and Mason used the Knopp–Mason canonical factorization [KM09] $f=f_{0}f_{1}$ (over $\mathbf{C}$) of a parabolic generalized modular function $f$ on a congruence subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$, where $f_{0}$ is a parabolic generalized modular function of a unitary character $\chi$, while $f_{1}$ is a parabolic generalized modular function without zeros or poles on the extended upper-half plane [KM03a]. Combining to their earlier method from [KM08], they thus proved that the unbounded denominators conjecture for the case of parabolic GMF is equivalent to the algebraicity of the first “few” Fourier coefficients of the component $f_{1}$ in the canonical factorization of $f$. As an application they proved Mason’s unbounded denominators conjecture for the case of a cuspidal parabolic GMF of weight 0 on a congruence group.
@@
In the case of two-dimensional representations of $\Gamma(2)$, Mason’s conjecture was settled by Franc and Mason [Mas12, FM14], and extended further by Franc, Gannon and Mason [FGM18] to the stronger sense of only requiring the $p$-adic boundedness of the coefficients for a density one set of primes $p$. Their proof relies on the special property that the power series in $\mathbf{Q}\llbracket x\rrbracket$ which arise in their context are hypergeometric functions. It is conceivable that the algebraicity part (over $\mathbf{Q}(x)$, respectively over the ring of classical modular forms) in Theorems 7.2.1 and 7.3.3 could likewise hold under a similar loosening of the integrality condition; but our proof does not imply this. On the other hand, for representations of dimension $n\geq 3$, it is plain that the congruence property ceases to hold as in [FM14] if we relax $\mathbf{Z}\llbracket q\rrbracket$ to $\mathbf{Z}[1/S]\llbracket q\rrbracket$. Another example where hypergeometric functions arise (this time for three-dimensional representations of $\mathrm{PSL}_{2}(\mathbf{Z})$) appears in the work of Franc–Mason [FM16a] and Marks [Mar15], and was employed by [FM16b] to derive certain cases of the original unbounded denominators conjecture.
@@ 7.4.2. Logarithmic vector-valued modular forms
If one drops the semisimplicity stipulation on $\rho\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)$ in the definition of a vector-valued modular form, the resulting structure has been named a logarithmic vector-valued modular form by Knopp and Mason [KM11]. They also do arise in conformal field theories, termed logarithmic (in place of rational). See, for example, Fuchs–Schweigert [FS19]. The components of a weight zero logarithmic vector-valued modular form with bounded denominators can now be classical (congruence) modular forms of higher weight, and so certainly transcendental over $\mathbf{C}(\lambda)$ (see Remark 7.2.3). In Question 7.4.5 below, we give an extension of the unbounded denominators problem over to the logarithmic setting. It remains outside the scope of our method as far as we could see. Before stating this question, we recall some basic facts concerning quasi-modular forms. Recall that the ring of quasi-modular forms $\widetilde{M}(\Gamma)$ for $\Gamma\subset\mathrm{PSL}_{2}(\mathbf{Z})$ may be identified with the ring generated by $E_{2}$ over the ring of classical holomorphic modular forms $M(\Gamma)$ of integral weight for $\Gamma$ [Zag08, Prop 20(ii)], and by [Zag08, Prop 20(i)] it is stable under the operator
@@
The ring of weakly holomorphic quasi-modular forms $\widetilde{M}^{!}(\Gamma)$ for $\Gamma\subset\mathrm{PSL}_{2}(\mathbf{Z})$ is the ring $\widetilde{M}(\Gamma)[1/\Delta]$.
@@
Question 7.4.5.
If a component $F_{j}(\tau)$ of a logarithmic vector-valued modular form for $\mathrm{PSL}_{2}(\mathbf{Z})$ has a $\mathbf{Z}\llbracket q\rrbracket$ Fourier expansion, does $F_{j}(\tau)$ belong to the ring of weakly holomorphic quasi-modular forms $\widetilde{M}^{!}(\Gamma)$ for some congruence subgroup $\Gamma\subset\mathrm{PSL}_{2}(\mathbf{Z})$?
@@
We now turn to some basic examples hinting towards a positive answer to question 7.4.5. Complementing Example 7.2.3 is the $\lambda$-pullback of the complete elliptic integral of the second kind:
@@
clearly a component of a logarithmic vector-valued modular form on $\mathrm{PSL}_{2}(\mathbf{Z})$, whose $q$-expansion is in $\mathbf{Z}\llbracket q\rrbracket$. But one can indeed verify that
$\frac{2}{\pi}E(\lambda(q))=\frac{\vartheta_{3}\vartheta^{4}_{4}+4\theta\vartheta_{3}}{\vartheta^{3}_{3}}$
@@
exhibiting this $\mathbf{Z}\llbracket q\rrbracket$ power series as an element of $\widetilde{M}^{!}(\Gamma)$ for the congruence subgroup $\Gamma=\Gamma(12)$, in accordance with Question 7.4.5.
@@ 7.4.7. Beyond $\mathrm{SL}_{2}(\mathbf{Z})$
Secondly, the mapping class groups $\Gamma_{g,n}=\mathrm{Mod}(S_{g,n})$ in signatures $(g,n)$ other than $(1,1),(1,0)$ or $(0,4)$ that we have implicitly been limiting to. Recall that $\Gamma_{1,1}\cong\Gamma_{1,0}=\mathrm{Mod}(\mathbf{T}^{2})=\mathrm{SL}_{2}(\mathbf{Z})$ and $\Gamma_{0,4}\cong\mathrm{PSL}_{2}(\mathbf{Z})\ltimes(\mathbf{Z}/2\times\mathbf{Z}/2)$, and correspondingly the discussion in the rational conformal field theory under § 7.1 has been for the $1$-loop partition function with a complex torus ($g=1$) as the worldsheet [Gan06]. In a more recent research stream in two-dimensional conformal field theory, a higher genus extension of Zhu’s modularity theorem would associate to any holomorphic vertex operator algebra a Teichmüller modular form (as defined in [Ich94]) in every signature $(g,n)$; this is a section of a tensor power $\lambda^{\otimes k}$ of the Hodge bundle over $\overline{\mathcal{M}_{g,n}}$. One could ask about extending the cruder algebraicity proviso of our Theorem 7.3.3 over to the more general setting of a component of a vector-valued Teichmüller modular form that has an appropriate integrality property.
7.4.8. The Ramanujan–Petersson question for cuspidal vector-valued modular forms
In the bulk of our paper, our theorems were stated for modular forms of an arbitrary integral [^5] They even hold for the modular forms of half-integral weight, upon multiplying by a weight-$1/2$ theta function. weight, but the path to arithmetic algebraization methods was always through a straightforward reduction to weight 0. This is because it did not make a difference in our integrality questions as to whether or not the forms in question were cuspidal. At the same time, all our theorems can be equivalently (without loss of generality) stated for the cuspidal forms. This offers a common global framework for unifying our Theorem 7.3.3 with the classical Ramanujan–Petersson conjectures.
Consider a representation $\rho:\mathrm{PSL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$. For each integer $k\in\mathbf{Z}$, denote by $S_{k}(\rho)$ the $\mathbf{Q}$-vector space of the weight $k$ vector-valued modular form whose multiplier system is $\rho$ and whose $n$ components all belong to $q^{1/N}\mathbf{Q}\llbracket q^{1/N}\rrbracket$ for some (unspecified) $N\in\mathbf{N}_{>0}$. We can focus our question on the $q$-expansion coefficients of an arbitrary $q\,\mathbf{Q}\llbracket q\rrbracket$ component $f(q)=\sum_{l=1}^{\infty}a_{l}q^{l}$ of an element of $S_{k}(\rho)$. The growth of those coefficients acquires a global, multicolored meaning upon completing $\mathbf{Q}\hookrightarrow\mathbf{Q}_{v}$ at the different places $v\in M_{\mathbf{Q}}$ of $\mathbf{Q}$. Denote by $|\cdot|_{v}$ the usual absolute value normalized by $|e|=e$, if $v=\infty$, and by $|p|_{p}=1/p$, if $v$ is the $p$-adic place. We can define a set $\Sigma(\rho)\subset M_{\mathbf{Q}}$ of the deficient places for the multiplier system $\rho$ by declaring $v\in\Sigma(\rho)$ if and only if there exists such a Fourier series component $f(q)=\sum_{l=1}^{\infty}a_{l}q^{l}\in q\,\mathbf{Q}\llbracket q\rrbracket$, for some weight $k\in\mathbf{Z}$ and some cuspidal vector-valued modular form $F\in S_{k}(\rho)$ to the multiplier system $\rho$ that has $f$ as one of its $n$ component functions, such that
[…]
The conjunction of Shimura’s integrality theorem [Shi59, § 8] and Deligne’s resolution [Del74, Théorème 8.2] of the Ramanujan–Petersson conjecture are exactly packaged together into the statement that $\Sigma(\rho)=\emptyset$ (no deficiency places) for the case that the representation $\rho:\mathrm{PSL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$ has kernel a congruence subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$. On the other hand, our proof of Theorem 7.3.3 established that $\Sigma(\rho)\neq\emptyset$ (and, more precisely, contains at least one non-archimedean place) in every other case, to wit: whenever the kernel $\ker(\rho)$ of the multiplier representation $\rho:\mathrm{PSL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$ is not a congruence subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$. As far as we are aware, it is an open question whether there exists any $\rho$ with $\infty\in\Sigma(\rho)$, but also whether there exists any $\rho$ with $\infty\notin\Sigma(\rho)$ but yet with $\ker(\rho)$ not a congruence subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$. We note however Selberg’s result [Sel65, § 2] that at least $|a_{l}|_{\infty}=O\left(l^{\frac{k}{2}-\frac{1}{5}}\right)$ (an improvement over Hecke’s trivial $\ll l^{k/2}$ bound, based on the Rankin–Selberg method) is in place for a completely arbitrary $\rho:\mathrm{PSL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$. Another question is whether the deficient places set $\Sigma(\rho)$ is always finite for a given multiplier representation $\rho:\mathrm{PSL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$. Eisenstein’s theorem [BG06, §11.4] guarantees that this is so in the case that the representation $\rho$ has a finite image.
@@ 7.4.9. Algebraic fundamental groups
Let $N\in\mathbf{N}_{>0}$, let $K/\mathbf{Q}(\mu_{N})$ be a finite extension, and let $\pi:\mathcal{X}(N)\to\mathrm{Spec}\,{O_{K}}$ (“connected Néron model”) be the connected component containing the cusp $\infty$ in the smooth part of the minimal regular model of $X(N)$ over $\mathrm{Spec}\,{O_{K}}$. Thus the cusp $\infty$ extends to a morphism $\varepsilon:\mathrm{Spec}\,{O_{K}}\to\mathcal{X}(N)$.
@@
This follows rather formally by the argument of [Iha94, § 4 on page 252] and also [Iha94, proof of Theorem 1 loc.cit. on pages 248–249], upon replacing Ihara’s function field $k(t)$ by the modular function field $K(X(N))$ and Ihara’s formal power series ring $\mathfrak{O}\llbracket t\rrbracket$ by $O_{K}\llbracket\lambda(\tau/N)/16\rrbracket$, taking account of Remark 6.3.1, and on using our Theorem 1.0.1 in place of Harbater’s arithmetic rationality input [Iha94, Claim 1A on page 248]. ∎
@@
by the Lerch–Chowla–Selberg formula making Bost’s capacitary condition [Bos99, Corollary 1.3] apply, and this is the isolated minimum value of the Faltings height across all elliptic curves. In practice this means that this complex torus has a “large” univalent complex-analytic uniformization (in the sense of conformal size from the origin $\infty$] and potential theory), sufficient to place this particular case of Theorem 7.4.10 within the framework of arithmetic rationality — as opposed to algebraicity or holonomicity — theorems [Bos99, BCL09] on the algebraic curve $X(N)$. Can such an approach be continued to all $N$?
@@ 8. Acknowledgments
We would like to thank Yves André for a number of insightful remarks on the first version of this manuscript, leading in particular to § 2.4.1 as an alternative and simplified approach to Theorem 2.0.1. We would also like to thank Michael Barz, Jean-Benoît Bost, François Charles, Pierre Deligne, Cameron Franc, Igor Frenkel, Javier Fresán, Jayce Getz, Kenz Kallal, Mark Kisin, Geoffrey Mason, Peter Sarnak, Alex Smith, Richard Taylor, John Voight, and Wadim Zudilin for useful remarks, suggestions, and corrections.